OpenStudy (anonymous):

Find the minimum product of two numbers who difference is 12. What are the two numbers? (How did you get it?)

OpenStudy (anonymous):

3 and 4 (I looked at the multiplication table)

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

is it -11

OpenStudy (anonymous):

How Matricked?

OpenStudy (anonymous):

because the answer is -36 and the two numbers are -6, 6

OpenStudy (anonymous):

but you said product, right? So I don't get it....Sorry

OpenStudy (anonymous):

product is multiply

OpenStudy (anonymous):

I know and that's why I'm asking :(

OpenStudy (anonymous):

oops my ans is incorrect yes it may be -36

OpenStudy (anonymous):

Now I am totally confused...........Sorry

OpenStudy (anonymous):

but how? and why?

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

yes. The question asks for: FInd the minimum product of two numbers whose difference is 12 and then answer is -36. Also, what are the two numbers: -6, 6

OpenStudy (anonymous):

6, -6

OpenStudy (anonymous):

How, A.M \(\ge \) GM

OpenStudy (anonymous):

A.M? GM?

OpenStudy (anonymous):

arithmetic mean and geometric mean respectively.

OpenStudy (anonymous):

put one number x, the other is x + 12 minimize \[x(x+12)=x^2+12x\] vertex is \[(-6,-36)\]

OpenStudy (anonymous):

ohh so how does solve -6, 6 if the vertex is -6,-36?

OpenStudy (anonymous):

satellite satellite !!! :)

OpenStudy (anonymous):

find vertex via \[x=-\frac{b}{2a}\]

OpenStudy (anonymous):

two numbers are x and x+ 12 we find x = -6 and therefore we know other one is -6 + 12 = 6

OpenStudy (anonymous):

ohh... how did you come up with x+12?

OpenStudy (anonymous):

ok lets imagine one number is 5 and the difference is 12 what is the other number?

OpenStudy (anonymous):

huh? so it is 7? o.o

OpenStudy (anonymous):

i get 17 because 17 - 5 = 12 how did i get it? added 12 to 5

OpenStudy (anonymous):

Btw A.M vs G.M generally works for non-negative numbers but (some what surprisingly) works here

OpenStudy (anonymous):

ohh.. i get it alright foodformath :)

OpenStudy (anonymous):

put one number x, the other is x + 12 minimize Let Product p = x(x+12)=x2+12x use Maxima Minima method to find the least product. differentiate the function p' = 2x + 12 for maxima or minima p' =0 => 2x + 12 =0 => x=-6 ............. the first no. => x+12 = +6 .... the second no. as f'' = 2 ........ a positive no. so the function has a minima at x= -6 and not a maxima.